# Cooper quartets designing in multi-terminal superconducting devices

### v1

@@
Dissipationless transport.— It becomes natural at this point to search for a direct consequence of having a quartet ground state and a smoking gun observable that can unambiguously witness its presence. We first look at the equilibrium properties and dissipationless transport in a three-terminal structure constituted by a common lead at the left, relabeled lead 0, and two additional leads 1 and 2 each coupled to the corresponding quantum dot, as schemitized in Fig. 2a), and study the supercurrents in the system. Proximity induced local pairing in quantum dot $i=1,2$ can originate either from lead 0 or from the corresponding lead, 1 or 2, at the right, biased at phase $\varphi_{i}$ with respect to lead 0, as schematized in Fig. 2 a). We assume different local rates at the left $\gamma_{S,l}$ and at the right $\gamma_{S,r}$ leads and equal for the two dots. This way, local pairing in dot $i$ becomes controlled by $\gamma_{S,i}=\gamma_{S,l}+\gamma_{S,r}e^{i\varphi_{i}}$. In turn, non-local pairing $\gamma_{S}^{\prime}$ can only result from Cooper pair splitting from lead 0 and it is not sensitive to any phase differences. The Hamiltonian projected on the low energy manifold reads
$h$ $=$ $\frac{\gamma_{S}^{2}}{W}(2+\cos(\varphi_{1})+\cos(\varphi_{2}))$ (8)
@@
-- $2\delta\epsilon(|\bar{0}\rangle\langle\bar{0}|-|\bar{4e}\rangle\langle\bar{4e}|)+\Gamma|\bar{4e}\rangle\langle\bar{0}|+{\rm H.c.},$
@@
that generalizes the coupling between vacuum and four-electron states to a complex one that, for $\gamma_{S,l}=\gamma_{S,r}=\gamma_{S}$, reads
$\Gamma(\varphi_{1},\varphi_{2})=-\frac{2(\gamma_{S}^{\prime})^{2}}{U+W}+\frac{\gamma^{2}_{S}}{W}(1+e^{i\varphi_{1}})(1+e^{i\varphi_{2}}).$ (9)
@@
Its argument $\Theta={\rm arg}\left[\Gamma\right]$ represents the phase of the superposition between vacuum and four-electron state in the ground state, $|\Psi_{\rm GS}\rangle=\frac{1}{\sqrt{2}}(|\bar{0}\rangle-e^{i\Theta}|\bar{4e}\rangle)$, and can be tuned through the phases $\varphi_{1}$ and $\varphi_{2}$.
@@
Figure 3: a) Current $I_{1}(\varphi_{1},\varphi_{2}=0)$ and b) $I_{2}(\varphi_{1},\varphi_{2}=0)$ as a function of the phase $\varphi_{1}$ and the detuning $\delta\epsilon$, obtained by keeping $\varphi_{2}=0$. c) Cuts of $I_{1}(\varphi_{1},\varphi_{2}=0)$ in a) at three different values of $\delta\epsilon$ shown in the legend of d). d) Cuts of $I_{2}(\varphi_{1},\varphi_{2}=0)$ at the values of $\delta\epsilon$ specified in the legend.
The most intuitive property a quartet ground state is expected to show is to support a two-Cooper pair current between externally attached superconducting leads. At an ordinary Josephson junction the tunneling of a Cooper pair amounts to a $2\pi$ slip in the phase difference at the junction, so that the coherent tunneling of two Cooper pairs yields a $\pi$-periodic Josephson current. The ground state energy is shown in Fig. 2b) and it clearly shows two minima at $\varphi_{1}=\varphi_{2}\simeq\pm\pi/2$. The equality does not hold in a strict way as a consequence of weak but finite single Cooper pair contribution in the lowest energy states $|\bar{0}\rangle$ and $|\bar{4e}\rangle$. Associated to the quartet ground state, we expect the system to support a quartet current, so that we join contact 1 and 2 in a unique drain by setting $\varphi_{1}=\varphi_{2}=\varphi$ and study the total current $I=I_{1}+I_{2}$. The ground-state current through dot $i$ features two contributions: an ordinary Josephson term describing current between leads $i$ and 0, and a second term that depends on both $\varphi_{1}$ and $\varphi_{2}$,
$I_{i}=-\frac{2e}{\hbar}\frac{\gamma_{S}^{2}}{W}\sin(\varphi_{i})-\frac{2e}{\hbar}\frac{\partial}{\partial\varphi_{i}}\left|\Gamma(\varphi_{1},\varphi_{2})\right|.$ (10)
@@
It is clear that a $\pi$-periodic current can only arise approximately, in that the contribution coming from $|\Gamma|$ must cancel the ordinary $2\pi$-periodic pair contribution of the first term. Assuming $W=-U-\delta W$, $\delta W\ll U$, $\gamma_{S}^{\prime}=\gamma_{S}$, and expanding the current at lowest orders in $\delta W/U$ we obtain
@@
The cancellation of the first harmonic is possible through a destructive interference effect originating from the opposite sign in the expression of $\Gamma$ between the phase-independent term describing non-local Cooper pair splitting and the phase-dependent term resulting from local Cooper pair tunneling. The result can be understood as follows: if we could ideally select only Cooper pair splitting at the left lead, by setting the local contribution to zero $\gamma_{S,l}=0$, and only direct Cooper pair tunneling at the right—as the structure with two terminals naturally selects—the only possible process would be a two-Cooper pair current, as shown in Fig. 2c) for comparison and explained in Note (1). For finite $\gamma_{S,l}=\gamma_{S,r}=\gamma_{S}$, we can obtain it approximately by promoting the relative importance of the non-local process with respect to the local one in the expression of $\Gamma$ by requiring $|2(\gamma_{S}^{\prime})^{2}/(U+W)|\gg|\gamma_{S}^{2}/W|$ so to effectively suppress the single Cooper pair current. The resulting current is shown in Fig. 2c) and shows a dominant $-\sin(2\varphi)$ component. Importantly, we obtain such a condition without recurrence to any flux interferometric methods Douçot and Ioffe (2012).
@@
The interacting nature of the quartet correlator suggests us that another possible clear signature of the quartet ground state can appear in the non-local response of the system. Indeed, a given phase difference established between two given terminals affects the current through the third terminal in a non-trivial way. If we close lead 2 on lead 0 and enclose a flux $\Phi_{x}$ we effectively pin the phase $\varphi_{2}=2e\Phi_{x}/\hbar$, so that we can selectively choose an horizontal cut in the energy landscape Fig. 2b) and study the associated currents in terminal 1 and 2. The resulting $I_{1}$ is shown in Fig. 3a) where application of a phase bias $\varphi_{1}$ between terminal 1 and 0 at constant $\varphi_{2}=0$ yields a sinusoidal current $I_{1}(\varphi_{1})$ away from the quartet resonance condition that sharply evolves in an opposite current at resonance $\delta\epsilon=0$, in an effective $\pi$-junction behavior. The latter is visualized in Fig. 3c) for three values of $\delta\epsilon$. Even more counter-intuitive is the current $I_{2}(\varphi_{1})$ that flows in the ring at zero external flux. $I_{2}$ activates only at resonance and shows a finite signal concord to $I_{1}$, as shown in Fig. 3b) and d).
Conclusions.— We have demonstrated how appropriate tuning of electron-electron interaction in a double quantum dot system opens the possibility to design quartet superconductivity. The system presented shows a great host of counter-intuitive properties and shall open unexplored paths in the simulation of correlated states of matter that hardly occur in nature, from charge-$4e$ superconductivity to more exotic phases based on novel electron complexes. Once a solid quartet ground state is stabilized a first phenomenological expectation is flux quantization in units of $h/4e$. To this end, contacts $S_{1}$ and $S_{2}$ need to be joined in a right superconducting lead, with care to not generate non-local Andreev reflection at the right lead, as explained in Note (1). Upon closing the circuit on a superconducting loop enclosing a flux $\Phi_{x}$, a charge $4e$ current appear as the $\pi$-periodic current in Fig. 2c), with $\varphi=2e\Phi_{x}/\hbar$. A consequent very appealing application is to design a parity-protected qubit, where two-Cooper pair transport yields quantum states with a definite parity of the number of Cooper pairs Blatter et al. (2001); Ioffe et al. (2002); Ioffe and Feigel’man (2002); Brooks et al. (2013); Douçot and Ioffe (2012). The possibility to experimentally achieve the isolation of a quartet ground state relies on the ability to engineer an attractive density-density interaction, that can be mediated by a second double quantum dot that is capacitively coupled to the system Hamo et al. (2016), or through a transmission line resonator Delbecq et al. (2013). Our results stimulate further research to consolidate the platform and holds promises for the design and simulation of highly correlated phases of matter based on complex electron aggregates.
@@
Acknowledgment.—We acknowledge the EU’s Horizon 2020 Research and Innovation Framework Programme under Grant No. 964398 (SUPERGATE), No. 101057977 (SPECTRUM), and the PNRR MUR project PE0000023-NQSTI for partial financial support. A.B. acknowledges the Royal Society through the International Exchanges between the UK and Italy (Grants No. IEC R2 192166.)

### v2

@@ III Correlated Andreev matter
To generalize the description we focus on a three-terminal structure constituted by a common lead at the left, relabeled lead 0, and two additional leads 1 and 2 each coupled to the corresponding quantum dot, as schematized in Fig. 2a), and study the supercurrents in the system. A proper description requires enlarging the basis and separately consider four two-electron states together with the vacuum $|0\rangle$ and the four-electron state $|4e\rangle$, so that the full basis reads $\{|0\rangle,|{\uparrow\downarrow}\rangle_{1}$, $|{\uparrow\downarrow}\rangle_{2}$, $|{\uparrow}\rangle_{1}|{\downarrow}\rangle_{2}$, $|{\downarrow}\rangle_{1}|{\uparrow}\rangle_{2},|4e\rangle\}$. Proximity-induced local pairing in quantum dot $i=1,2$ can originate either from lead 0 or from the corresponding $i$-lead at the right which is biased at phase $\varphi_{i}$ with respect to the reference lead 0, as schematized in Fig. 2a). We assume different local rates at the left $\gamma_{S,l}$ and at the right $\gamma_{S,r}$ leads, equal for the two dots. This way, local pairing in dot $i$ becomes controlled by $\gamma_{S,i}=\gamma_{S,l}+\gamma_{S,r}e^{i\varphi_{i}}$. In turn, non-local pairing $\gamma_{S}^{\prime}$ can only result from Cooper pair splitting from lead 0 and it is not sensitive to any phase differences.
The low energy states are modified by the complex rates and now read
@@
$|\bar{0}\rangle$ $=$ $|0\rangle+\sum_{i}\frac{\gamma_{S,i}}{2W-2\delta\epsilon}|{\uparrow\downarrow}\rangle_{i}+\frac{\sqrt{2}\gamma_{S}^{\prime}}{U+W-2\delta\epsilon}|\phi^{-}_{2e}\rangle,$
$|\bar{4e}\rangle$ $=$ $|4e\rangle+\frac{\gamma^{*}_{S,2}|{\uparrow\downarrow}\rangle_{1}+\gamma^{*}_{S,1}|{\uparrow\downarrow}\rangle_{2}}{2W-2\delta\epsilon}-\frac{\sqrt{2}\gamma_{S}^{\prime}}{U+W-2\delta\epsilon}|\phi^{-}_{2e}\rangle,$
and the Hamiltonian projected on the low-energy manifold reads
[…]
@@
with $𝟙\openone$ a $2\times 2$ identity in the low energy manifold and $\Gamma$ a generalized complex matrix element between vacuum and four-electron states, that at $\delta\epsilon=0$ reads
$\Gamma(\varphi_{1},\varphi_{2})=-\frac{2(\gamma_{S}^{\prime})^{2}}{U+W}+\frac{1}{W}(\gamma_{S,l}+\gamma_{S,r}e^{i\varphi_{1}})(\gamma_{S,l}+\gamma_{S,r}e^{i\varphi_{2}}).$ (13)
@@
Its argument $\Theta(\varphi_{1},\varphi_{2})={\rm arg}\left[\Gamma\right]$ represents the phase of the superposition between vacuum and four-electron state in the ground state, $|\Psi_{\rm GS}\rangle=\frac{1}{\sqrt{2}}(|\bar{0}\rangle-e^{i\Theta}|\bar{4e}\rangle)$, and can be tuned through the phases $\varphi_{1}$ and $\varphi_{2}$.
The ground state energy as a function of the two phase differences $\varphi_{1},\varphi_{2}$ is shown in Fig. 2b) and it clearly shows two minima at $\varphi_{1}=\varphi_{2}\simeq\pm\pi/2$ along the $\varphi_{1}=\varphi_{2}$ line, and a saddle point at $\varphi_{1}=\varphi_{2}=0$, manifesting the strongly correlated character of the ground state. In particular, we notice that if the leads $S_{1}$ and $S_{2}$ are left floating, so that the system adjusts in one minimum of the ground state energy, the latter breaks time-reversal symmetry.
The Hamiltonian Eq. (12) well captures also the first excited state, whose exact landscape is shown in Fig. 2c) together with the ground state at zero detuning $\delta\epsilon$. The system realizes a quartet Andreev qubit that represents an instance of correlated Andreev matter beyond the family of multi-terminal Josephson junctions Riwar et al. (2016); Teshler et al. (2023); Ohnmacht et al. (2024); Zalom et al. (2024), and possibly featuring topological properties that will be discussed in future works.
IV Dissipationless transport
@@
Figure 3: a) Current $I_{1}(\varphi_{1},\varphi_{2}=0)$ and b) $I_{2}(\varphi_{1},\varphi_{2}=0)$ as a function of the phase $\varphi_{1}$ and the detuning $\delta\epsilon$, obtained by keeping $\varphi_{2}=0$. c) Cuts of $I_{1}(\varphi_{1},\varphi_{2}=0)$ in a) at three different values of $\delta\epsilon$ shown in the legend of d). d) Cuts of $I_{2}(\varphi_{1},\varphi_{2}=0)$ at the values of $\delta\epsilon$ specified in the legend. The parameters for the plots are the same as in Fig. 2.
We now analyze the dissipationless transport properties of the system. The ground-state current through dot $i$ features two contributions: an ordinary Josephson term describing current between leads $i$ and 0, and a second term that depends on both $\varphi_{1}$ and $\varphi_{2}$,
$I_{i}=-\frac{2e}{\hbar}\frac{\gamma_{S,l}\gamma_{S,r}}{W}\sin(\varphi_{i})-\frac{2e}{\hbar}\frac{\partial}{\partial\varphi_{i}}\left|\Gamma(\varphi_{1},\varphi_{2})\right|.$ (14)
We first look for a two-Cooper pair current in a two-terminal configuration by joining contact 1 and 2 in a unique drain by setting $\varphi_{1}=\varphi_{2}=\varphi$ and study the total current $I=I_{1}+I_{2}$. An ideal two-Cooper pair current flows in the case $\gamma_{S,l}=0$. This is easily understood as follows. Starting in the quartet ground state a Cooper pair coming from the left lead can only split between the two dots, by virtue of the condition $\gamma_{S,l}=0$. Such a state is current-blocked by the superconducting gap in the leads 1 and 2, that have no single-particle states in the spectrum (especially in the $\Delta\to\infty$ limit). A second Cooper pair coming from the left lead can again only split and result in a four-electron state in the double-dot. The latter is now unblocked and gives rise to an ideal two-Cooper pair current, as shown in Fig. 2d). This intuitive picture is confirmed by the form of the ground state energy that, setting $\gamma_{S,r}=\gamma_{S}$, reads
$E_{\rm GS}=-E_{0}\sqrt{1+\tau\sin^{2}(\varphi)},$ (15)
with $E_{0}=\left|-\frac{2(\gamma_{S}^{\prime})^{2}}{U+W}+\frac{\gamma_{S}^{2}}{W}\right|$ and $\tau=\frac{8(\gamma_{S}\gamma_{S}^{\prime})^{2}}{E_{0}^{2}W(U+W)}$, that is manifestly $\pi$-periodic. In addition, for $\tau>0$ the minima of the ground state energy Eq. (15) are at $\varphi=\pm\pi/2$, so that the current at small finite bias flows in the direction opposite to the phase bias, in an effective $\pi$-junction behavior, as shown in Fig. 2d).
Figure 4: a) Schematics of a possible implementation of the system. The double quantum dot is formed on a carbon nanotube in tunnel contact with superconducting leads $S_{0}$, $S_{1}$, and $S_{2}$. Gates $V_{g1}$ and $V_{g2}$ allow for alignment of the dots levels. A second carbon nanotube, on the right side is suspended between contacts $S_{1}$ and $S_{2}$ but not in tunnel contact. Its vibrating lowest energy flexural mode couples to the charge on the quantum dots and provided a mechanism for strong interdot density-density interaction. b) Modified setup that allows the pinning of the phase $\varphi_{2}=2e\Phi_{x}/h$ and the resulting current $I_{1}$ shown in Fig. 3. c) Quartet correlator $Q$ away from resonance, with $\delta\epsilon/\gamma_{S}=0.1$ and d) $\delta\epsilon/\gamma_{S}=1$, with $\gamma_{S}^{\prime}=\gamma_{S}$.
@@
For the realistic case of $\gamma_{S,l}=\gamma_{S,r}=\gamma_{S}$, the different terms compete and a $\pi$-periodic current can only arise approximately, in that the contribution coming from $|\Gamma|$ must cancel the ordinary $2\pi$-periodic pair contribution of the first term Eq. (14). This can be achieved by promoting the relative importance of the non-local process with respect to the local one in the expression of $\Gamma$ by requiring $|2(\gamma_{S}^{\prime})^{2}/(U+W)|\gg|\gamma_{S}^{2}/W|$. Crucial is the opposite sign between the terms that originates from the singlet expression of the non-local Cooper pair tunneling and yields a destructive interference effect. By setting $W=-U-\delta W$, $\delta W\ll U$, $\gamma_{S}^{\prime}=\gamma_{S}$, and expanding the current at lowest orders in $\delta W/U$ we obtain
@@
The resulting current is shown in Fig. 2d) and shows a dominant $-\sin(2\varphi)$ component. Importantly, we obtain such a condition without recurrence to any flux interferometric methods Douçot and Ioffe (2012).
V Non-local phase response
@@
The interacting nature of the quartet correlator suggests us that another possible clear signature of the quartet ground state can appear in the non-local response of the system. A non-local response appears by noticing that a given phase difference established between two given terminals, say 0 and 2, affects the current through the third terminal in a non-trivial way. If we close lead 2 on lead 0 and enclose a flux $\Phi_{x}$ (as schematized in Fig. 4b)) we effectively pin the phase $\varphi_{2}=2e\Phi_{x}/\hbar$, so that we can selectively choose an horizontal cut in the energy landscape Fig. 2c). The resulting current $I_{1}$ is shown in Fig. 3a) as a function of the phase $\varphi_{1}$ and the detuning $\delta\epsilon$ and for three specifics values of $\delta\epsilon$ in Fig. 3c): the application of a phase bias $\varphi_{1}$ between terminal 1 and 0 at constant $\varphi_{2}=0$ yields a sinusoidal current $I_{1}(\varphi_{1})$, away from the quartet resonance condition, that sharply evolves in a current of opposite sign at resonance $\delta\epsilon=0$, in a sort of effective $\pi$-junction behaviour. At the same time, at resonance a current $I_{2}(\varphi_{1})$ starts to flow in the ring at zero external flux, with the same negative sign as $I_{1}$, as shown in Fig. 3b) and d).
This behavior constitutes a clear signature of the quartet Andreev process for this simple setup, that becomes active only at resonance, and it is understood as follows: away from the resonance the ground state has a weak quartet component and an ordinary dissipationless current flows from terminal 2 to terminal 0, following the phase bias. As resonance is approached, the system develops a ground state with a strong quartet component, that necessarily yields currents in both terminal 1 and 2, that goes in the opposite direction to the phase bias, as dictated by the interference between the phase dependent and the phase independent term that constitute $\Gamma$.
It is worth to remind that the emergence of a $\pi$-junction behavior typically requires engineering complex structures, that involve ferromagnetic systems Ryazanov et al. (2001), unconventional order parameters Tsuei and Kirtley (2000), non-equilibrium effects Baselmans et al. (1999), or weak Kondo correlations Cleuziou et al. (2006). Furthermore, the $I_{2}$ current is anomalous in that it flows in a zero phase bias condition in absence of any time-reversal breaking mechanism other than the phase $\varphi_{1}$. Typically anomalous Josephson effect requires strong spin-orbit interaction and a Zeeman field Szombati et al. (2016), conditions that are not verified in our case.
VI Possible experimental setup
Here, we provide a possible experimental implementation of a system that could host the attractive double quantum dot and several other possibility have been explored in literature Hamo et al. (2016); Delbecq et al. (2013); Bhattacharya et al. (2021); Vigneau et al. (2022); Moser et al. (2013). The system is shown in Fig. 4a) and it is composed by two carbon nanotubes: the left one can be placed to be in tunnel contact with the three superconducting contacts $S_{0}$, $S_{1}$, and $S_{2}$, with $S_{0}$ at the center, and the right one suspended between contacts $S_{1}$ and $S_{2}$, but not galvanically coupled, and in close proximity with the left one. This way, the contacts define two dots and the central one can inject Cooper pairs locally and non-locally. The charge on the two dots couples to the charge on the suspended nanotube via Coulomb interaction in the vacuum. The latter is not screened and the coupling is thus given by the long range Coulomb interaction $\propto 1/r^{2}$. The bending of the suspended nanotube depends on the total charge on the double quantum dots and the minimal Hamiltonian consisting of a single nanotube flexural mode $a$ reads
$H=\omega_{0}a^{\dagger}a+g_{0}(a+a^{\dagger})(n_{1}+n_{2}).$ (17)
with $a$ denoting he bosonic annihilation operator of the flexural mode. The Hamiltonian is quadratic in the bosonic modes, so that it can be rewritten as $H=\omega_{0}A^{\dagger}A-(g_{0}^{2}/\omega_{0})(n_{1}+n_{2})^{2}$, where the new displaced bosonic mode reads $A=a+(g_{0}/\omega_{0})(n_{1}+n_{2})$. The ground state will be given by the vacuum of $A$, so that the correction of the energy results in the effective interaction for the two dots
$H_{\rm int}=U(n_{1\uparrow}n_{1\downarrow}+n_{2\uparrow}n_{2\downarrow})+Wn_{1}n_{2},$ (18)
with $U=U_{0}-\frac{2g_{0}^{2}}{\omega_{0}}$ and $W=W_{0}-\frac{2g_{0}^{2}}{\omega_{0}}$, with $U_{0}>0$ and $W_{0}>0$ the bare onsite repulsion and interdot density-density interaction strengths. The crucial point is that the bare frequency of the flexural phonon can be very low compared to the coupling to the electronic density Mahan and Jeon (2004), so that a strong renormalization of the bare interaction is possible. For bare energy scales $U_{0}\simeq 1~{}{\rm THz}$, $\omega_{0}\simeq 1~{}{\rm MHz}$, and $g_{0}\simeq 1~{}{\rm GHz}$, values of the coupling constant $0.350.25<g_{0}^{2}/(U_{0}\omega_{0})<$ 0.35 can be realized Bhattacharya et al. (2021); Vigneau et al. (2022); Moser et al. (2013).
VII Conclusions
We have demonstrated how appropriate tuning of electron-electron interaction in a double quantum dot system opens the possibility to engineer a quartet superconductor. The system presented shows a great host of counter-intuitive properties and opens unexplored paths in the simulation of correlated states of matter and the study of interactions effects superconductors. As a further development in the context of transport in hybrid normal-quantum-dot-superconducting devices we foresee the possibility of a quartet Andreev reflection, something that may make the quartet ground state nature evident and that will be addressed in future works.
The possibility to experimentally achieve the isolation of a quartet ground state relies on the ability to engineer an attractive density-density interaction. The latter has been mediated by a second double quantum dot capacitively coupled to the system Hamo et al. (2016), or through a transmission line resonator Delbecq et al. (2013). In addition, we have presented a simple possible setup based on a suspended carbon nanotube as a possible effective mediator of the attractive interdot interaction.
The quartet resonance is fragile to detuning $\delta\epsilon$, as shown in Fig. 1d), but also to temperature, in that the correlator is washed away for temperature $T$ larger than $|\Gamma|$, that separates the ground state $|\phi^{-}_{Q}\rangle$ from the excited state $|\phi^{+}_{Q}\rangle$. Accounting for the latter results in $Q\simeq-(1-e^{-|\Gamma|/T})/2$. Nevertheless, if the hierarchy of scales $T,\delta\epsilon<|\Gamma|<\gamma_{S},\gamma_{S}^{\prime}\ll U,|W|<\Delta$, is respected the picture holds. Assuming $\gamma_{S},\gamma_{S}^{\prime}\simeq 20~{}{\rm GHz}$ Biercuk et al. (2005); Hofstetter et al. (2009); Herrmann et al. (2010), we expect that temperatures of the order of $25$ mK should be enough to stabilize a quartet ground state. Furthermore, relaxing the infinite gap condition does not change drastically the picture, so long as the entire double-dot spectrum falls within the superconducting gap $\Delta$.
In the finalization of the work, we became aware of recent studies concerning quartet superconductivity in a two-orbital Hubbard model with similar results Soldini et al. (2024).
@@ VII Conclusions
Acknowledgment.— We acknowledge useful discussions with W. Belzig, L. Amico. We acknowledge the EU’s Horizon 2020 Research and Innovation Framework Programme under Grant No. 964398 (SUPERGATE), No. 101057977 (SPECTRUM), and the PNRR MUR project PE0000023-NQSTI for partial financial support. The work has been also supported by the Fondazione Cariplo on the project N. 2023-2594. A.B. acknowledges MUR-PRIN 2022 - Grant No. 2022B9P8LN-(PE3)-Project NEThEQS “Non-equilibrium coherent thermal effects in quantum systems” in PNRR Mission 4-Component 2-Investment 1.1 “Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN)” funded by the European Union-Next Generation EU, the Royal Society through the International Exchanges between the UK and Italy (Grants No. IEC R2 192166) and CNR project QTHERMONANO.
Appendix A Correlator
When the Wick theorem applies, the expectation value of any product of operators on the ground state can be decomposed in all possible two-operator contractions. In the case of quartet superconductivity, the Wick theorem does not apply and the correlator itself has been defined as cleaned by the two-point contractions, so that when the state is a Slater determinant the correlator is zero. In the definition we have omitted the equal-spin two-point contractions, that are zero in our case.
In general the Hamiltonian Eq. (5) gives rise to a complex matrix representation. However, for the case of a single superconducting contact or for no phase differences between the terminals in the case of a multi-terminal structure, it is possible to choose a gauge in which the Hamiltonian is real and so are the eigenstates. Assuming a generic state
$|\psi\rangle$ $=$ $\cos(\theta_{1})|0\rangle+\sin(\theta_{1})\cos(\theta_{2})|\phi^{+}_{2e}\rangle$ (19)
$+$ $\sin(\theta_{1})\sin(\theta_{2})\cos(\theta_{3})|\phi^{-}_{2e}\rangle$
$+$ $\sin(\theta_{1})\sin(\theta_{2})\sin(\theta_{3})|4e\rangle,$
The maximum of $|Q|$ is obtained for the values $\theta_{1}=\pi/4$ and $\theta_{2}=\theta_{3}=\pm\pi/2$, corresponding to the states $|\phi^{\pm}_{Q}\rangle$ for which the correlator takes the values $Q=\pm 1/2$. Numerical checks on eight-component even parity states with real coefficients confirm the bounds $-1/2\leq Q\leq 1/2$.
The value of the correlator is therefore very significant of the quartet content of the ground state and for completeness in Fig. 4c) and d) we show its value slightly away from exact resonance, for $\delta\epsilon/\gamma_{S}=0.1,1$ and for $\gamma_{S}^{\prime}=\gamma_{S}$.